OpenStudy (anonymous):

I am using Pearson's correlation coefficient to find the correlation of my data. The x values clearly increase while my y values decrease. Why is my r value positive?

OpenStudy (perl):

r shouldn't be positive, it should be close to -1. perhaps you are looking at r^2

OpenStudy (anonymous):

In my graphic calculator using LinReg r = -0.797 but when I manually do Pearson's correlation coefficient r is 1.

OpenStudy (perl):

then you have a mistake in your manual work. can i see the work you did and i will need the original data

OpenStudy (perl):

checking over, one minute

OpenStudy (perl):

that sum says you have to add Numerator: $$ \large {\sum(x-\bar x)(y- \bar y) \\ = (x_1 - \bar x)\cdot (y_1 - \bar y) + (x_2 - \bar x )(y_2 - \bar y) + ... (x_6 - \bar x) (y_6 - \bar y) } $$

OpenStudy (perl):

in the denominator you have $$ \large {\sum(x-\bar x)^2 = (x_1 - \bar x)^2 + (x_2 - \bar x )^2 + ... (x_6 - \bar x) ^2 \\ \sum(y-\bar y)^2 = (y_1 - \bar y)^2 + (y_2 - \bar y )^2 + ... (y_6 - \bar y) ^2 } $$

OpenStudy (anonymous):

Thank you! :)

OpenStudy (perl):

I modified the excel a bit, so its more clear

OpenStudy (perl):

if click on the cells, you can see the calculation

OpenStudy (anonymous):

Ok, thank you! :)